# Probing Langmuir monolayer self-assembly in condensed and collapsed phases: grazing incidence X-ray diffraction and X-ray standing waves studies

### v1

@@ 2.3 Grazing incidence diffraction
The time evolution of the monolayer in the collapse mode 2 is shown in Fig. 6. In this series of measurements, the monolayers underwent a rapid structural change compared to the measurement timescale. We measured both XSW and GID. To do this, we prepared several monolayer samples in collapse mode 2. In Fig. 6 we show GID data from two different samples. For both samples all controllable conditions were the same. On the first sample, we performed the GID measurements immediately after compression. The GID pattern measured within the first 2 hours is shown in Fig. 6(a), and the data collected between 3 and 11 hours after compression are shown in Fig. 6(b). On the second sample, we first performed the XSW measurements, which will be discussed in the next section, starting at 3 hours after compression we performed several GID measurements in between XSW scans. This is done in such order to analyze the dynamics of the Ce distribution during the structural change in collapse mode 2. The GID pattern collected between 3 and 5 hours after compression is shown in Fig. 6(c) and between 22 and 24 hours is shown in Fig. 6(d).
@@
As in all other measurements, the monolayer initially formed into a structure with the NN tilt packing mode [see Fig. 6(a)].However, as the measurements continued, the GID revealed the structural change. Fig. 6(b) shows a rich diffraction pattern, which we assume to be a superposition of two patterns: three lower intensity peaks of PHB located exactly at the same coordinates (within 1% accuracy) as the PHB peaks in series 2; and two brighter peaks similar to a typical picture of the NN tilt, but with coordinates different from those in all other series. Also, unlike all other measured NN tilt patterns, the peak above the horizon is to the right of the peak on the horizon. Furthermore, one can see that the bright peak above the horizon is slightly rounded. The roundness of the peaks is clearly observed in the GID pattern in Figs. 6(c), 6(d) Three rounded peaks, again at , characteristic of the PHB, are seen in Fig. 6(c). No other peaks are in this pattern. This measurement is taken 3.5 and 5 hours after compression of the monolayer. It is noteworthy that the GID pattern measured 24 hours after compression is similar [see Fig. 6(d)], indicating that the monolayer is relatively stable in collapse mode 2. The GID pattern is different in terms of the maximum height of the peaks, but there are still all three PHB peaks and they are also rounded.
@@ 2.4 X-ray standing waves
We applied X-ray standing wave (XSW) technique to monitor the changes in the location of cerium atoms during the self-assembling reorganization of the AA monolayer. In XSW studies at liquid interfaces, the intensity of secondary radiation (e.g., characteristic fluorescence), excited by the incident X-rays, is measured as a function of the incident angle in the subcritical region . The angular dependence of fluorescence intensity is known to be highly sensitive to the position of the atoms emitting secondary radiation signal, that allows to locate the atoms by the fitting the corresponding fluorescence curve. For quantitative analysis of XSW data the fluorescence intensity curves are typically simulated using the Parratt equations [25] or optical transfer matrices (see [26, 27] among others).
@@
The XSW measurements for the solid-phase monolayer were carried out 21 hours after the AA monolayer was compressed to mN/m. Fig. 7(a) shows the angular dependence of integrated intensity for Ce peak obtained in these measurements. From this figure, the intensity of the Ce-fluorescence increases dramatically from zero at and reaches the sharp maximum near the critical angle . This is a typical behavior, which is observed for the angular dependence of the fluorescence intensity from atoms distributed in the near-surface region [28]. Theoretical simulations have shown that there is a good agreement between the calculations and the experimental data, for the solid-phase monolayer, which can be obtained within the framework of a two-layer model: the Ce-free top layer (AA monolayer) and the Ce-containing bottom. Both layers were considered to be homogeneous. The theoretical Ce-fluorescence curve in Fig. 7(a) was calculated for the case when Ce atoms are present in a 5 Å thick layer under the top layer with a thickness of 27 Å. Based on these data we can suggest that Ce atoms formed a single layer underneath the AA monolayer, that is a quite reasonable arrangement of Ce atoms for a solid-phase monolayer.
@@
To understand the general trends in the changes in the location of cerium ions it is convenient to use the weight coefficient ratio , that allows to estimate the fractional amount of the cerium ions, located above the liquid surface. The theoretical angular dependencies of the fluorescence intensity in Fig. 7(c), 7(d) were calculated for the following values of this parameter: curve 2: , curve 3: , curve 4: and curve 5: . It should be taken into account that good agreement between the theoretical and the experimental curves could be obtained also by varying the thickness of the Ce distribution above the liquid surface. In such a case the changes in the angular dependence of Ce-fluorescence from curve 2 to curve 3 can be explained by the increase of the thickness of the Ce distribution above the liquid surface; whereas the changes in curves 3–5 clearly indicate that this parameter gradually decreases.
Thus, the XSW data obtained for collapse mode 2 provided a clear idea about the Ce-induced processes in the collapsed AA monolayer: In the initial stage (during the first 3 hours), the collapse events are governed by an increase in the thickness of the Ce distribution above the liquid surface and/or an increase in the fraction of Ce atoms located above the liquid surface. In the late stage (during further 19 hours of measurements) the reorganization processes in the AA film are reversed. Namely, the thickness of the Ce distribution above the liquid surface decreases and/or the fractional amount of Ce atoms located above the liquid surface decreases.
@@ 3 Discussion
Let us now look at all of the above observations from the perspective of the molecular organization of the monolayer. First, note that the arachidic acid monolayer formed on a Ce3+ solution is thermodynamically very different from the monolayer formed on pure water, as shown by the - isotherms in Fig. 1.
@@
The GID data provide information about the structure i.e. translational order of the monolayer. We performed GID measurements on the monolayers with gradually increasing structural complexity. First, we consider the monolayer in the solid phase. Compared to the control series with pure water subphase, the monolayer formed on a subphase with Ce3+ ions has new structural features. The Ce3+ ions form an inorganic superstructure underneath the monolayer: a periodically arranged inorganic layer with a unit cell. This is evidenced by the peaks in range A. These peaks are higher than those of the organic structure, with full width half maximum in the vertical direction Å-1, from which the vertical size of the structure Å can be estimated to be smaller than that of the arachidic acid hydrocarbon chain Å. The XSW data are insensitive to lateral order, but unlike the GID, they are sensitive to the position of the Ce layer in the vertical direction. The fluorescence intensity curve calculated for the Ce layer underneath the arachidic acid monolayer is in good agreement with the data in Fig. 7(a). This calculation is made for the 5 Å thick Ce layer. This is in agreement with the estimation using the GID. Taken together these, results imply the Ce layer is indeed underneath the organic monolayer, and it does form a laterally periodic inorganic superstructure. This is also consistent with the analysis in [10, 11]. Regarding the chemical composition of the inorganic superstructure, we assume that no nitrate ligands are present in the nearest environment of cerium ions in Ce(NO3)3 salt solution at the concentration used in our experiments. This assumption is based on the results reported in [29], where X-ray absorption spectroscopy is used to study the interaction of Ln3+ ions with nitrate ligands in aqueous solutions of Ln(NO3)3 salts. The formation of an inorganic superstructure has an effect on the structure of the monolayer itself. Compared to the monolayer on a pure water subphase, it is more densely packed with 2% less area occupied by a molecule. This is apparently realized by the NN tilt of the AA hydrocarbon chains.
@@
Another structural feature is revealed in the similarly prepared monolayer formed at the lower temperature C: the phase mode characterized by the PHB hydrocarbon chains ordering. At higher temperatures, the molecular chains are in the free rotatory state. The hydrocarbon chain have a flattened shape. At lower temperatures, the motion of the molecules is reduced and the shape of the chains begins to play a role. It is then possible to reduce the area occupied by a molecule by aligning the hydrocarbon chains of neighboring molecules. When the azimuth angle between adjacent chains is 90∘, it is a herringbone phase mode, and when it is around 40∘, it is a PHB. It is illustrative to mention the following. Consider a plot with and lattice parameters as axes. A point on this plot would represent a monolayer with certain lattice. Also consider the parameters , which are the projections of the unit cell onto the plane perpendicular to the molecular chains. In other words, the parameters represent the shortest distance between molecules and are therefore more relevant when considering lattice energy calculations. It is described in the literature [5] that monolayers of amphiphilic molecules in the solid phase lie on an arc in this plot. Fig. 8 shows the contour plot of the distribution of different studied monolayers, it is based on the data published in [22, 30, 31]. Moreover it is known [22, 32] that the upper left corner of this plot is associated with PHB. In Fig. 8 all the structures considered in our study [see also Table 2.3] are marked with triangles and their corresponding transverse parameters are marked with diamonds. The structures we observed in our study also obey this phenomenological rule: all the diamond markers in Fig. 8 are on this ark. Note that the solid phase monolayer formed at C is in the upper right corner of this plot. In addition, the PHB phase mode is confirmed by the presence of the (2/3 0) peak in the GID pattern. It should also be noted that the structure of the monolayer has changed. Immediately after compression, the monolayer had a structure similar to that of the monolayer at C. However, with time, a PHB phase mode appeared. The PHB mode coexisted with the NN tilted phase mode. In the PHB, the hydrocarbon chains are not tilted because all corresponding peaks are on the horizon. In Fig. 8 it is shown as follows: the triangular marker corresponding to this structure lies exactly on the diamond, as .
@@
It is quite the opposite for the type of structural ordering we refer to as collapse mode 2. The collapse mode 2 was observed at C. First, the BAM images of collapse mode 2 exhibit a mosaic-like structure [see Figs. 2(e), 2(f)]. Second, we were able to observe the GID pattern over all 24 hours of GID measurements that implies that the monolayer retains translational order. This order is maintained over 24 hours of measurements. According to our GID experiments, once the monolayer is beyond the collapse point, it initially retains the NN packing mode, further NN coexists with the PHB, and then only the PHB packing mode remains. See the evolution of the GID pattern in Fig. 6; note that the patterns in Figs. 6(a), 6(b) and in Figs. 6(c), 6(d) are measured on different samples, but all controllable parameters were the same. The PHB packing mode is maintained for an extended period of time once collapse mode 2 is reached. Although the patterns in Figs. 6(c), 6(d) can clearly be attributed to the PHB, since their coordinates measured at the horizon Å-1 match those in series 2 (cf. Fig. 3), the patterns have one unusual feature: the peaks are rounded. We qualitatively associate these roundings with the corrugation of the monolayer. Consider a monolayer with domains that are allowed to lean from being flat on the liquid surface. Unlike the usual GID pattern, which is formed by cylindrical averaging over all possible crystallite positions and thus has a rod-like shape, the GID pattern for the corrugated monolayer is formed not only by this, but also by averaging over all possible orientations of the domains with respect to the surface. Hence the rounding of the peaks. Such peaks must lie on the circle with the origin of the reciprocal space in its center, which is exactly what we observed. However, these peaks are not closed to complete a ring as in 3D powder diffraction. This is because the greater the slope of a domain from the flat position, the less likely that position is in the ensemble.
@@
It is of immediate interest to have a numerical characterization of such a structure. A simple analysis of peak positions and widths provides the orientational and translational structure, and also an estimate of the average size of a single crystal in the 2D powder, if the resolution limit of the experimental setup allows. In most practical cases such an analysis is sufficient. However, when analyzing partially disordered states such as collapse mode 2, more information about the structure is required. This may include the parameters of the angular distribution of the domains (flat single crystal regions of the corrugated film) or whether the film is partially or fully corrugated. Such an analysis requires numerical simulations based on the physical principles of X-ray diffraction. This approach is much less commonly reported. In this regard, we refer to the works in [33, 34], where the effect of the structure of the fatty acid molecules is carefully considered to calculate line cuts in the 2D GID patterns. In [35], an approach is proposed to calculate the full 2D GID pattern from a polycrystalline monolayer. Two kinds of structural imperfections are taken into account in [35]: thermal vibrations of the molecules leading to Deby-Waller type decay of the diffraction amplitude, and the average size of a single crystal in the monolayer.
@@
Using the new approach, we performed the simulations of the GID patterns. For the simulations, we assumed that the orientation angle of the domains (the angle between the normal to the surface and the normal to the domain) is normally distributed and that the domains are sufficiently large so that the resulting diffuse scattering is averaged incoherently. We also assumed that there is no preferred azimuthal orientation of the domain, i.e. there is no mosaicity. For simplicity, we also assumed that each domain is a flat single crystal. For the forward simulations we consider the following parameters: average domain size, lattice parameters, tilt angle, effective thickness of the monolayer, standard deviation of the orientation angle and the ratio of the corrugated area to the total area of the monolayer. In the context of this research, we focus on the latter two parameters: and . The simulations were fit to the experimental data by optimizing the model parameters using the Levenberg-Marquardt algorithm [37]. The resulting best-fit theoretical GID patterns are shown in Fig. 9. First, note that peak rounding is adequately reproduced in the simulations. The simulation in Fig. 9(b) is done for the normally distributed orientation angle, and Fig. 9(d) is a weighted sum between the normally distributed slope of the domain and a simulation for the flat domains. In both the experiment [see Fig. 9(c)] and the simulation [see Fig. 9(d)] one can see the faint straight peak near the brighter rounded one. This is consistent with the observations in XSW where the Ce3+ ions first upwards above the liquid surface and then the process was reversed: a part of the monolayer became flat again. The best fit standard deviation of the domain slope distribution for the first GID pattern [see Fig. 9(b)] is . For the second pattern [in Fig. 9(d)] is . The best fit for the latter considers a mixture of flat domains and the rest being normally distributed. This observation is also in qualitative agreement with the BAM measurements.

### v2

@@ 2.3 Grazing incidence diffraction
The time evolution of the monolayer in the collapse mode 2 is shown in Fig. 6. In this series of measurements, the monolayers underwent a rapid structural change compared to the measurement timescale. We measured both XSW and GID. To do this, we prepared several monolayer samples in collapse mode 2. In Fig. 6 we show GID data from two different samples. For both samples all controllable conditions were the same. On the first sample, we performed the GID measurements just after compression. The GID pattern measured within the first 2 hours is shown in Fig. 6(a), and the data collected between 3 and 11 hours after compression are shown in Fig. 6(b). On the second sample, we first performed the XSW measurements, which will be discussed in the next section, starting at 3 hours after compression we performed several GID measurements in between XSW scans. This experimental strategy allowed us to analyze the dynamics of the Ce distribution during the structural change in collapse mode 2. The GID pattern collected between 3 and 5 hours after compression is shown in Fig. 6(c) and between 22 and 24 hours is shown in Fig. 6(d).
@@
As in all other measurements, the monolayer initially formed into a structure with the NN tilt packing mode [see Fig. 6(a)]. However, as the measurements continued, the GID revealed the structural change. Fig. 6(b) shows a rich diffraction pattern, which we assume to be a superposition of two patterns: three lower intensity peaks of PHB located exactly at the same coordinates (within 1% accuracy) as the PHB peaks in series 2; and two brighter peaks similar to a typical picture of the NN tilt, but with coordinates different from those in all other series. Also, unlike all other measured NN tilt patterns, the peak above the horizon is to the right of the peak on the horizon. Furthermore, one can see that the bright peak above the horizon is slightly rounded.
In the pattern measured on another sample [see Fig. 6(c)], the roundness of the peaks is clearly visible. There are no other peaks associated with the NN tilt mode in this pattern. The peaks are on circles centered at the origin of the reciprocal space. The pattern remained unchanged while the surface pressure was kept constant by shrinking the surface area. This continued until the minimum position of the trough barrier was reached 5 hours after the initial compression. After that, the GID pattern changed again. The pattern in Fig. 6(c), obtained 24 hours after compression, has a subtle difference. The peaks in this pattern appear to be partly rounded and partly straight. We assume that this GID pattern is the incoherent sum of two patterns: one with rounded peaks and another with straight peaks. However, to confirm this assumption, numerical simulations of evanescent X-ray diffraction are required. This will be considered further, but first we would like to point out another noteworthy observation of series 4. Namely, there are no inorganic superstructure peaks [see Fig. 3(e)]. This immediately raises the question of the localization of Ce in the collapse mode 2. This question can be answered with the help of X-ray standing waves.
@@ 2.4 X-ray standing waves
We applied X-ray standing wave (XSW) technique to monitor the changes in the location of cerium atoms during the self-assembling reorganization of the AA monolayer. In XSW studies at liquid interfaces, the intensity of secondary radiation (e.g., characteristic fluorescence), excited by the incident X-rays, is measured as a function of the incident angle in the subcritical region . The angular dependence of fluorescence intensity is known to be highly sensitive to the position of the atoms emitting secondary radiation signal, that allows to locate the atoms by the fitting the corresponding fluorescence curve. For quantitative analysis of XSW data the fluorescence intensity curves are usually simulated using the Parratt equations [25] or optical transfer matrices (see [26, 27] among others).
@@
The XSW measurements for the solid-phase monolayer were carried out 21 hours after the AA monolayer was compressed to mN/m. Fig. 7(a) shows the angular dependence of integrated intensity for Ce peak obtained in these measurements. From this figure, the intensity of the Ce-fluorescence increases dramatically from zero at and reaches the sharp maximum near the critical angle . This is a typical behavior, which is observed for the angular dependence of the fluorescence intensity from atoms distributed in the near-surface region [28]. Theoretical simulations have shown that there is a good agreement between the calculations and the experimental data, for the solid-phase monolayer, which can be obtained within the frame of a two-layer model: the Ce-free top layer (AA monolayer) and the Ce-containing bottom layer. Both layers were considered to be homogeneous. The theoretical Ce-fluorescence curve in Fig. 7(a) was calculated for the case when Ce atoms are present in a 5 Å thick layer under the top layer with a thickness of 27 Å. Based on these data we can suggest that Ce atoms formed a single layer underneath the AA monolayer, that is a quite reasonable arrangement of Ce atoms for a solid-phase monolayer.
@@
To understand the general trends in the changes in the location of cerium ions it is convenient to use the weight coefficient ratio , that allows to estimate the fractional amount of the cerium ions, located above the liquid surface. The theoretical angular dependencies of the fluorescence yield in Fig. 7 (c),(d) were calculated for the following values of this parameter: for curve 2, for curve 3, for curve 4 and for curve 5.
According to these results the distinct trend in changes of XSW data from curve 1 to curve 3 can be explained by an enhancement of the fractional amount of the Ce atoms located above the liquid surface. Importantly, these XSW measurements were performed, while keeping the surface pressure constant, i.e. under continuous compression. Curves 4 and 5 were recorded at markedly different conditions, namely when the minimal area was attained and the surface pressure fell down. Not surprisingly, the changes in the shape of these Ce-fluorescence curves represent completely different processes – the decrease of the fractional amount of the Ce atoms located above the liquid surface.
It should be taken into account that good agreement between the theoretical and experimental Ce-fluorescence curves could be obtained also by varying the thickness of the Ce distribution above the liquid surface. In such a case the changes in the angular dependence of Ce-fluorescence from curve 2 to curve 3 can be described by the increase of the thickness of the Ce distribution above the liquid surface; whereas the changes of Ce-fluorescence curves 3–5 clearly indicate that this parameter gradually decreases.
Thus, XSW data provided a clear idea about the Ce-induced processes in collapsed AA monolayer at C. During first 5 hours, when the AA monolayer was held at constant surface pressure, a pronounced changes in monolayer reorganization can be attributed to an increase in the thickness of the Ce distribution above the liquid surface or/and an enhancement of the fractional amount of the Ce atoms located above the liquid surface. During further 19 hours of measurements, when the through barrier reached its uttermost position, the events occurring in collapsed AA film were reversed. In particular, the thickness of the Ce distribution above the liquid surface decreases or/and the fractional amount of the Ce atoms located above the liquid surface reduces.
@@ 3 Discussion
Let us now look at all of the above observations from the perspective of the molecular organization of the monolayer. First, note that the AA monolayer formed on a Ce3+ solution is thermodynamically very different from the monolayer formed on pure water, as shown by the - isotherms in Fig. 1.
@@
We performed GID measurements on the monolayers with gradually increasing structural complexity. First, we consider the monolayer in the solid phase. Compared to the control series with pure water subphase, the monolayer formed on a subphase with Ce3+ ions has new structural features. The Ce3+ ions form an inorganic superstructure underneath the monolayer: a periodically arranged inorganic layer with a unit cell. This is evidenced by the peaks in range A. These peaks are higher than those of the organic structure, with full width half maximum in the vertical direction Å-1, from which the vertical size of the structure Å can be estimated to be smaller than that of the AA hydrocarbon chain Å. The XSW data are insensitive to lateral order, but unlike the GID, they are sensitive to the position of the Ce layer in the vertical direction. The fluorescence intensity curve calculated for the Ce layer underneath the AA monolayer is in good agreement with the data in Fig. 7(a). This calculation is made for the 5 Å thick Ce layer. This is in agreement with the estimation using the GID. Taken together these results imply that the Ce layer is indeed underneath the organic monolayer, and it does form a laterally periodic inorganic superstructure. This is also consistent with the analysis in [10, 11]. Regarding the chemical composition of the inorganic superstructure, we assume that no nitrate ligands are present in the nearest environment of cerium ions in Ce(NO3)3 salt solution at the concentration used in our experiments. This assumption is based on the results reported in [29], where X-ray absorption spectroscopy is used to study the interaction of Ln3+ ions with nitrate ligands in aqueous solutions of Ln(NO3)3 salts. The formation of an inorganic superstructure has an effect on the structure of the monolayer itself. Compared to the monolayer on a pure water subphase, it is more densely packed with 2% less area occupied by a molecule. This is apparently realized by the NN tilt of the AA hydrocarbon chains.
@@
Another structural feature is revealed in the similarly prepared monolayer formed at the lower temperature C: the packing mode characterized by the PHB hydrocarbon chains ordering. At higher temperatures, the molecular chains are in the free rotatory state. The hydrocarbon chain have a flattened shape. At lower temperatures, the motion of the molecules is reduced and the shape of the chains begins to play a role. It is then possible to reduce the area occupied by a molecule by aligning the hydrocarbon chains of neighboring molecules. When the azimuth angle between adjacent chains is 90∘, it is a herringbone phase mode, and when it is around 40∘, it is a PHB. It is illustrative to mention the following. Consider a plot with and lattice parameters as axes. A point on this plot would represent a monolayer with certain lattice. Also consider the parameters , which are the projections of the unit cell onto the plane perpendicular to the molecular chains. In other words, the parameters represent the shortest distance between molecules and are therefore more relevant when considering lattice energy calculations. According to the literature [5] that monolayers of amphiphilic molecules in the solid phase lie on an arc in this plot. Fig. 8 shows the contour plot of the distribution of different studied monolayers, it is based on the data published in [22, 30, 31]. Moreover it is known [22, 32] that the upper left corner of this plot is associated with PHB. In Fig. 8 all the structures considered in our study [see also Table 2.3] are marked with triangles and their corresponding transverse parameters are marked with diamonds. The structures we observed in our study also obey this phenomenological rule: all the diamond markers in Fig. 8 are on this ark. Note that the solid phase monolayer formed at C is in the upper right corner of this plot. In addition, the PHB phase mode is confirmed by the presence of the (2/3 0) peak in the GID pattern. It should also be noted that the structure of the monolayer has changed. Immediately after compression, the monolayer had a structure similar to that of the monolayer at C. However, with time, a PHB phase mode appeared. The PHB mode coexisted with the NN tilted phase mode. In the PHB, the hydrocarbon chains are not tilted because all corresponding peaks are on the horizon. In Fig. 8 it is shown as follows: the triangular marker corresponding to this structure lies exactly on the diamond, as .
@@
It is quite the opposite for the type of structural ordering we refer to as collapse mode 2. The collapse mode 2 was observed at C. First, the BAM images of collapse mode 2 exhibit a mosaic-like structure [see Figs. 2(e), 2(f)]. Second, we were able to observe the GID pattern over all 24 hours of GID measurements that implies that the monolayer retains translational order. According to our GID experiments, once the monolayer is beyond the collapse point, it initially retains the NN packing mode, further NN coexists with the PHB, and then only the PHB packing mode remains. See the evolution of the GID pattern in Fig. 6; note that the patterns in Figs. 6(a), 6(b) and in Figs. 6(c), 6(d) are measured on different samples, but all controllable parameters were the same. The PHB packing mode is maintained for an extended period of time once collapse mode 2 is reached. Although the patterns in Figs. 6(c), 6(d) can clearly be attributed to the PHB, since their coordinates measured at the horizon Å-1 match those in series 2 (cf. Fig. 3), the patterns have one unusual feature: the peaks are rounded. We qualitatively associate these roundings with the corrugation of the monolayer. Consider a monolayer with domains that are allowed to lean from being horizontal on the liquid surface. Unlike the usual GID pattern, which is formed by cylindrical averaging over all possible crystallite positions and thus has a rod-like shape, the GID pattern for the corrugated monolayer is formed not only by this, but also by averaging over all possible orientations of the domains with respect to the surface. Hence the rounding of the peaks. Such peaks must lie on the circle with the origin of the reciprocal space in its center, which is exactly what we observed. However, these peaks are not closed to complete a ring as in 3D powder diffraction. This is because the greater the slope of a domain from the horizontal position, the less likely that position is in the ensemble.
@@
It is of immediate interest to have a numerical characterization of such a structure. A simple analysis of peak positions and widths provides the orientational and translational structure, and also an estimate of the average size of a single crystal in the 2D powder, if the resolution limit of the experimental setup allows. In most practical cases, such an analysis is sufficient. However, when analyzing partially disordered states such as collapse mode 2, more information about the structure is required. This may include the parameters of the angular distribution of the domains (flat single crystal regions of the corrugated film) or whether the film is partially or fully corrugated. Such an analysis requires numerical simulations based on the physical principles of X-ray diffraction. This approach is much less commonly reported. In this regard, we refer to the works in [33, 34], where the effect of the structure of the fatty acid molecules is carefully considered to calculate line cuts in the 3D GID patterns. In [35], an approach is proposed to calculate the full 2D GID pattern from a polycrystalline monolayer. Two kinds of structural imperfections are taken into account in [35]: thermal vibrations of the molecules leading to Deby-Waller type decay of the diffraction amplitude, and the average size of a single crystal in the monolayer.
@@
Using the new approach, we performed the numerical simulations to reproduce the GID patterns presumably associated with corrugated structures of collapse mode 2. As a model of a corrugated structure, we assumed the monolayer to be an ensemble of flat crystallite domains with randomly distributed slope. Here, the slope is the angle between the normal to the surface and the normal to the domain. Furthermore, we assume that the slope is normally distributed. We also assume that the domains are sufficiently large so that the resulting diffuse scattering is incoherently averaged. Finally, there is no preferred azimuthal orientation of the domain in our model. For the simulation we consider the following parameters: average domain size, lattice parameters, tilt angle, effective thickness of the monolayer and standard deviation of the slope . The simulations were fit to the experimental data by optimizing the model parameters using the Levenberg-Marquardt algorithm [37].
First, we consider the GID pattern measured 5 hours after compression in Fig. 6 (c) [also shown in Fig. 9 (a) for reference]. The best-fit simulation is shown in Fig. 9 (b). Note that peak rounding is adequately reproduced in the simulation. In other words, the corrugated monolayer model can quantitatively reproduce the experiment. On average, the slope of the domain is (domain is oriented horizontally on the surface). This is due to the position of the intensity maximum of the rounded peaks: exactly on the horizon. The best fit standard deviation of the domain slope distribution [for the simulation in Fig. 9(b)] is . This is to be interpreted as follows: approximately 70% of the domains are inclined with a slope less then .
The GID pattern measured 24 hours after compression is different [see the same data in Fig. 6 (d) and in Fig. 9 (c)]. The peaks in this pattern are partially rounded and partially straight as for the typical GID pattern of the flat monolayer. To simulate this GID pattern, we simply assumed that there are two coexisting phase modes: a corrugated and a flat one. We assumed that the diffraction on each part gives a non-coherent contribution to the final GID pattern. To account for this numerically, we introduced the ratio parameter into the model. The parameter describes the ratio of the corrugated area to the total area of the monolayer.
The best-fit simulation for this model, with . and , is shown in Fig. 9 (d). Again, the experimental data are reproduced: one can see rounded peaks together with faint straight peaks in Fig. 9 (d). The best-fit parameters can be interpreted as follows. At this stage, 63% of the monolayer is flat, as in the usual case. The other 37% of the monolayer is corrugated, but with a higher slope () than in the former case (). To avoid confusion, we note that the parameter represents the standard deviation of the slope, and not the slope itself. Such a mixed configuration of the monolayer (corrugated and flat) is likely related to the fact that 5 hours after compression, the barrier of the trough reached its minimum position and the surface pressure of the sample was no longer maintained. Remarkably, the monolayer compressed beyond the collapse point maintained its structural ordering even 24 hours after compression. On the contrary, part of the monolayer became flat again. It is difficult to say with the data we have whether this flattening was due to relaxation of the monolayer or some other unknown self-assembly process, and this is an interesting topic for further study.
These findings are confirmed by our XSW studies. As discussed above the results obtained in XSW measurements implies that during first 5 hours after the compression beyond the collapse point (while keeping the surface pressure constant) the collapse events in AA film manifested themselves in an increase in the thickness of the Ce distribution above the liquid surface or/and an enhancement of the fractional amount of the Ce atoms located above the liquid surface. After that, when the area decreased to minimum value and the surface pressure fell down, the thickness of the Ce distribution above the liquid surface decreases or/and the fractional amount of the Ce atoms located above the liquid surface reduces.
A final point to be addressed regarding collapse mode 2 is the fact that no inorganic superstructure of Ce3+ ions was observed in the GID data. This observation is not consistent with the XSW measurements, which prove the localization of Ce3+ ions beneath the AA monolayer. There are two possible explanations. The first is that the Ce layer is amorphous in collapse mode 2, so there is no diffraction. And the second is that the diffraction is attenuated by the distortions introduced into the Ce layer by the corrugated AA monolayer. In this case, one must explain why the diffraction from the AA structure is detectable while the diffraction from Ce is not. This may be due to the difference in effective size and shape of the scattering objects: AA molecules are much larger than the thickness of the Ce layer, and the same vertical distortions may affect the diffraction from these two structures differently.